HUBBLE SPACE TELESCOPE DAILY REPORT #1035 PERIOD COVERED: 0800 EST Daily Status Report as of 351/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of 5254 (GHRS SMOV Monitoring) The High Resolution Spectrograph was turned on at intervals of three hours, as part of a prerequisite to turning on the high voltage. A total of 15 observations were made, received and displayed. They all were reported to look normal. 1.2 Completed FOS/RD 4803 (SMOV FOS/RD Internal Operations) The Faint Object Spectrograph was used to make a set of dark calibration exposures and calibration exposure observations to verify the science data interfaces and proper functioning of the FOS (Red Side). These images will determine if the detector's electronics are working correctly, and the wavelength calibration verifies that the FOS optical path is still intact. Two observations were received in real time, and both images were reported to look fine. 1.3 Completed FOS/RD 5243 (FOS Location of Spectra and Aperture: SMOV Tests) The Faint Object Spectrograph (Red Side) was used to determine the coarse location for the 0.3 arcsecond aperture and the Y-location of the four spectral grating settings. A total of ten observations were received and displayed. All observations were reported to be normal. 1.4 Completed Two Sets of GHRS/2 4780 (SMOV GHRS Detector Calibrations) The High Resolution Spectrograph was used to make routine calibration observations as part of a standard detector test. All the exposures were internal images and did not require any special pointing. A total of four images were received and displayed in the OSS. They all were reported to look fine. 1.5 Completed GHRS/2 4757 (GHRS SMOV Light Path Verification) The High Resolution Spectrograph was used in the OBSERVE Mode, and standard calibration exposures were made using the Spectral Calibration Lamp #2. This activity will validate the instrument's optical path prior to any attempt to view an external target. One observation was received and displayed in the OSS. It was reported to look okay. 1.7 Completed FOC/96 5077 (FOC/96 SMOV Reverification Proposal) The FOC high voltage was turned on for the first time after the servicing mission, and four internal exposures were taken. Three images were flat fields and one was a dark exposure. These images will be compared to the pre-mission images to verify that the FOC is still in good health. All four images were received and displayed. They were reported to look okay. 1.8 Completed FOC/96 4679 (FOC SMOV Internal Flat Fields) The Faint Object Camera was used in parallel with other observations to obtain a series of internal flat field images. The data are required for determining the overall sensitivity and large scale characteristics of the FOC detector. Five internal flat observations were received and displayed in the OSS. They were reported to look okay. 1.9 Completed FOS 4802 (FOS SMOV Engineering Proposal) The Faint Object Spectrograph was used to make dark and calibration lamp exposures to verify the science data interfaces and the proper functioning of the FOS. The high voltage was on and reading 22.84 Kvolts. The Vacuum gauge read 0.316E-03 mT. Two images were received and displayed. The dark was reported to look okay, and the wavelength calibration was also as expected. No anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: No guide star acquisitions were taken during this period. 2.2 FHST Updates: Scheduled : 2 Successful : 2 2.3 Significant HSTARs: None 2.4 Operations Note: The SMOV planning activity and SMOV daily operations are on-going, with the STScI creating the SMSs that produce most of the HST commanding for this period. Both FOS Red Side and FOC/f96 began their initial operations, and were executed with no operational implications. GHRS calibration and verification activities continue. Gyros and magnetometers are performing nominally. Bias updates for the three gyro configuration are still converging. The Four Gyro FRR was conducted after successful VEST testing. This configuration is planned to be implemented today when gyros 1 and 2 are powered down. EPS status continues to be nominal, with average maximum battery SOC remaining at approximately 500 Ahrs and average SA output at 154 amps. Wing slewing is occurring nominally. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete S/C 4711 (WFPC2 Decontamination and KSPOT Check) 3.2 Complete WFPC2 4707 (WFPC2 Calibration System Flat Field Check) 3.3 Complete S/C 5241 (Guide Star Acquisition Verification) 3.4 Complete WFPC 2 4721 (WFPC2 Decontamination and TEC Turn-On ) 3.5 Complete S/C 5254 (GHRS SMOV Monitoring) 3.6 Complete WFPC2 5252 (WFPC2 AFM Turn-On And check-out) 3.7 Complete Two Sets of FGS 5425 (SMOV K-Factor Verification Test/Trans Scans For FGS 1, 2,and 3) 3.8 Complete Two Sets of WFPC2 4709 (WFPC2 Internal Monitor) 3.10 Complete Two Sets of WFPC2 4265 (WFPC2 Aperture Location and Internal Pickoff Mirror Adjust) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-